OpenStudy (anonymous):

i need to solve \[\cos \left( \frac{ \pi }{ 6 }+x \right)-\cos \left( \frac{ \pi }{ 6 }-x \right)=1\]

OpenStudy (anonymous):

Might help to use some trig identities. Give me a second.

OpenStudy (anonymous):

\[ \cos(x+y) = \cos(x)\cos(y) - \sin(x)\sin(y) \\ \]

OpenStudy (anonymous):

In which case you would treat \(x = \pi / 6\) and \(y = x\), etc.

OpenStudy (anonymous):

Forgive me for reusing that \(x\) variable.

OpenStudy (swissgirl):

lol You are forgiven :)

OpenStudy (swissgirl):

cos(x-y)=Cos(x)Cos(y)+Sin(x)Sin(y)

OpenStudy (anonymous):

thanks. i got it

OpenStudy (swissgirl):

What was your answer for x?

OpenStudy (anonymous):

\[\frac{ 3\pi }{ 2 }\]

OpenStudy (swissgirl):

Ya got that too :)

OpenStudy (anonymous):

good. yeah!